OpenStudy (anonymous):

Multiply. (x – 4)(x2 – 3x + 5)

mathslover (mathslover):

ok ! \[\large{(x-4)(x^2-3x-5)=x(x^2-3x-5)-4(x^2-3x-5)}\] can u solve this ?

OpenStudy (anonymous):

can you help me step by step ?

mathslover (mathslover):

ok ! see we have an identity : \(\huge{x(y-z)=xy-xz}\) or for trinomials we have : \[\huge{x(a-b-c)=xa-xb-xc}\] BINOMIALS * TRINOMIALS \[{(x-y)(a-b-c)=x(a-b-c)-y(a-b-c)=xa-xb-cx-ay+yb+yc}\] apply this identity here : given \(\huge{(x-4)(x^2-3x-5)}\) \[\huge{x(x^2-3x-5)-4(x^2-3x-5)}\] \[\large{x(x^2)-x(3x)-x(5)-4(x^2)-4(-3x)-4(-5)}\] \[\huge{x^3-3x^2-5x-4x^2+12x+20}\] \[\huge{x^3-3x^2-4x^2-5x+12x+20}\] U can simplify this further

OpenStudy (anonymous):

is it x3 + 5x2 + 5x – 20

mathslover (mathslover):

no @GhettoBaby \[\huge{x^3-3x^2-4x^2+12x-5x+20}\] \[\huge{x^3-7x^2+7x-20}\]

OpenStudy (anonymous):

you mean 17x-20

OpenStudy (anonymous):

@mathslover x3 – 7x2 + 17x – 20